A cliff is 22 meters high.You fell in the water 5 meters away from the hil.What is your initial horizontal velocity


1
Expert's answer
2021-11-11T09:32:13-0500

The time "t" of free fall (assuming no initial vertical velocity) can be found from kinematic equation:


"h = \dfrac{gt^2}{2}\\
t = \sqrt{\dfrac{2h}{g}}"

where "h = 22m" is the height of the cliff, and "g = 9.8m\/s^2" is the gravitational acceleration.

Since the horizontal velocity does not changes, its initial value is:


"v = \dfrac{d}{t} = d\sqrt{\dfrac{g}{2h}}"

where "d = 5m" is the horizontal range. Thus, obtain:


"v = d\sqrt{\dfrac{g}{2h}} =5m\cdot \sqrt{\dfrac{9.8m/s^2}{2\cdot 22m}} \approx 2.4m/s"

Answer. 2.4 m/s.